# Junyu Cao and Mohsen Bayati

### v1

@@
A Probabilistic Approach for Alignment with Human Comparisons
@@
A Probabilistic Approach for Alignment with Human Comparisons
@@
A growing trend involves integrating human knowledge into learning frameworks, leveraging subtle human feedback to refine AI models. Despite these advances, no comprehensive theoretical framework describing the specific conditions under which human comparisons improve the traditional supervised fine-tuning process has been developed. To bridge this gap, this paper studies the effective use of human comparisons to address limitations arising from noisy data and high-dimensional models. We propose a two-stage “Supervised Fine Tuning+Human Comparison” (SFT+HC) framework connecting machine learning with human feedback through a probabilistic bisection approach. The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SFT procedure, and then uses human comparisons to improve the model alignment. To examine the efficacy of the alignment phase, we introduce a novel concept termed the “label-noise-to-comparison-accuracy” (LNCA) ratio. This paper theoretically identifies the conditions under which the “SFT+HC” framework outperforms pure SFT approach, leveraging this ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity. We validate that the proposed conditions for the LNCA ratio are met in a case study conducted via an Amazon Mechanical Turk experiment.
@@
model alignment, supervised fine-tuning, probabilistic bisection, human-AI interaction.
@@ 1 Introduction
In light of the increasing availability of data, the demand for data-driven algorithmic solutions has surged. Machine learning has emerged as the standard technique for numerous tasks, including natural language processing and computer vision, among others. Its great importance has also been demonstrated across various domains of operations research.
Traditional methodologies have predominantly relied on supervised fine-tuning (SFT) [^1] Supervised fine-tuning is usually used in Large Language Models (LLM). We use supervised fine-tuning and supervised learning interchangeably in this paper. techniques by training on noisy-labeled data. A common practice in this context for decision-makers is to directly adopt the output of the best predictor identified by SFT, where the best predictor can be defined by the minimizer of the empirical risk, for example. However, some issues may arise in this approach. One significant challenge emerges from the prevalence of large noise, which can substantially hinder model performance. On the other hand, to mitigate the model misspecification issue, the growing trend toward increasingly complex models may lead to over-parameterization. Over-parameterization presents additional obstacles to the design of efficient and accurate AI systems. In instances where the model exhibits high complexity, such as when the feature dimension is comparable to or exceeds the number of data points, multiple distinct estimators may exhibit seemingly equivalent empirical performance on the training dataset but perform poorly on unseen data, a scenario exacerbated by the inherent under-specification of models (D’Amour et al. 2022). Thus, decision-makers must consider strategic data utilization and refine models to achieve a better performance under different environments.
@@
Recently, an ever-growing number of papers have incorporated human-knowledge into the learning framework (Christiano et al. 2017, Li et al. 2016). Fine-tuning with human feedback has shown a promising direction. For instance, in the design of chatbots, the performance can be adaptively and subsequently improved through the interactions with humans (Ziegler et al. 2019, Ouyang et al. 2022). Ouyang et al. (2022) focuses on fine-tuning approaches to aligning language models, where InstructGPT is fine-tuned with human feedback. In many cases, human evaluators possess domain expertise that goes beyond what can be encoded into a dataset or estimation method. Their deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al. 2020).
“Thinking is difficult, that’s why most people judge” – Carl Jung. Coming up with an answer is much more difficult than selecting a better answer from pair-wise comparisons. Zhou and Xu (2020) use pair-wise comparisons for story generation; Xu et al. (2023) collect human comparisons over two sampled LLM for the algorithm fine-tuning process. While prior research has showcased remarkable results when integrating human feedback across various applications, it has yet to establish a comprehensive theoretical framework detailing the specific conditions under which human comparison significantly augments the conventional SFT framework. To build on this research trajectory and fill this gap, our paper provides a theoretical framework for the strategic utilization of human comparisons. The primary focus is on addressing the inherent limitations posed by noisy data and the complexities of high-dimensional models. By actively and effectively leveraging human comparisons, we aim to navigate the intricacies of model alignment with the minimal labeling effort.
@@
To connect the machine learning framework with human feedback, we propose a two-stage tractable framework through a probabilistic method. In the first stage, noisy labeled data is fed into the SFT oracle to obtain low-dimensional representations. More specifically, the algorithm can use lasso to select important features from high dimensions or use matrix completion to learn low-rank representations. The second stage employs human comparisons to improve model alignment within this low-dimensional space. Our framework, which utilizes the random utility model to describe human’s comparison behavior, can incorporate the possibility of false selection in human comparisons. Human may provide incorrect responses, though this noise tends to be substantially lower than that present in labeled data. To accommodate this aspect, we develop a probabilistic bisection approach for the alignment phase. We call this two-stage process as “SFT+HC” framework.
The goal of our paper is to refine the model with a certain precision level and a confidence level using the minimum possible data. The rationale behind this framework design is rooted in the observation that some representation learning techniques, such as autoencoders or deep learning architectures (Zhou and Paffenroth 2017, Akrami et al. 2019), are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). These methods aim to capture underlying patterns and structures in the data. Recently, Tsou et al. (2023) show that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, offering competitive performance and, in some cases, even surpassing the results of full-model fine-tuning, where all weights are updated. In the second stage, we adopt human comparisons instead of human estimations because comparison is a much easier task. Our approach involves sequential design of comparisons and active data acquisition, with the objective of maximizing information and consequently minimizing sample complexity.
@@
To describe the effectiveness of the refinement phase through human comparison, we introduce a new concept called “label-noise-to-comparison-accuracy” (LNCA) ratio. We theoretically characterize the condition for the “SFT+HC” framework performing better than the pure SFT algorithm based on this LNCA ratio. In the large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
@@ 1.1 Our Contributions
Our contributions are summarized as follows:
  1. (I)
    Modeling: In a large-noise environment, machine learning models may exhibit fragility, due to the sensitivity to variations in input data and the dynamic nature of their surroundings. To address this challenge, we introduce a novel approach that leverages human comparisons to refine these models. To the best of our knowledge, our study represents the first theoretical attempt aimed at explaining the substantial enhancement of model performance under specific conditions through human comparisons. We present a novel two-stage framework based on a utility model for human comparisons. The model captures a natural phenomena that making distinctions between two options becomes increasingly challenging as they are more similar. Consequently, the accuracy of these comparisons may decrease as the model refines.
  2. (II)
    Theoretical Contributions:
    1. (a)
      Probabilistic bisection method: In the second phase, the decision-maker (DM) sequentially selects two options for human evaluators to compare. The fundamental question at this stage pertains to the efficient acquisition of valuable information from these human comparisons. To address this challenge, we introduce a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons. This method involves two types of actions: the ‘vertical move’ involves human evaluations of the same comparison. Once the algorithm identifies the superior choice with high probability, it proceeds with the ‘horizontal move,’ where the system selects the next pair for comparison. Our objective is to determine identify correct model parameters with a precision of $\epsilon$ and a confidence level of $1-\delta$. We establish the sample complexity required to achieve these precision and confidence levels.
    2. (b)
      Label-noise-to-comparison-accuracy ratio: We introduce a novel concept known as the ‘noise-to-human-knowledge ratio,’ which quantifies the relative scale between label noise and human-made comparison accuracy. In essence, when label noise is substantial, but humans can make selections with high accuracy, our two-stage framework outperforms pure Sequential Fine-Tuning (SFT). Specifically, when human selection accuracy deviates from 50%, the sample complexity for the refinement stage is $O(s\log(1/(\delta\varepsilon)))$. In cases where selection accuracy approaches 50% as the two options become indistinguishable from the true model, the sample complexity increases. We characterize the LNCA ratio conditions under which our SFT+HC framework outperforms pure SFT.
  3. (III)
    Numerical Performance: We evaluate the LNCA ratio through an empirical study on Amazon Mechanical Turk with a stylized task. The experiment result validates the condition under which “SFT+HC” performs better than pure SFT. Subsequently, we evaluated the performance of our proposed SFT+HC framework and compared it with the pure SFT approach across varying LNCA ratios. Our findings align with our theoretical results, revealing that the superiority of the SFT+HC framework becomes more pronounced under conditions of higher observational noise, higher accuracy of human comparisons, and lower dimensions of important features.
1.2 Literature Review
Alignment and Learning From Human Feedback. The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms towards the objectives and interests of humans. Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts, which was identified as underspecification (D’Amour et al. 2022). Predictors that perform similarly in the training dataset can exhibit significant variations when deployed in various domains. The presence of ambiguity in a model might result in instability and suboptimal performance in real-world scenarios. Recent papers have focused on aligning models with human intentions, including training simple robots in simulated environments and Atari games (Christiano et al. 2017, Ibarz et al. 2018), fine-tuning language models to summarize text fine-tuning (Stiennon et al. 2020), optimize dialogue (Jaques et al. 2019). It has been demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al. 2022, Glaese et al. 2022, Stiennon et al. 2020, Dwaracherla et al. 2024). Ouyang et al. (2022) use reinforcement learning from human feedback (Stiennon et al. 2020) to fine-tune GPT-3 to follow a broad class of written instructions. In particular, Zhou and Xu (2020) use pair-wise comparisons for story generation. They find that when comparing two natural language generation (NLG) models, instead of asking human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al. 2017), it is much easier for human annotators to directly compare one sample generated by the first model against another sample from the second model pairwisely and compute the win/loss rate. Zhu et al. (2023) provide reinforcement learning framework with human feedback, specifically, for function approximation for $K$-wise comparisons with policy learning as the target. Xu et al. (2023) propose a fine-tuning algorithm by inducing a complementary distribution over the two sampled LLM responses to model human comparison processes, in accordance with the Bradley-Terry model for pairwise comparisons. Different than those above-mentioned work, we provide a theoretical two-stage framework for integrating supervised learning with actively acquired human comparison labels, and characterize the conditions where the two-stage framework outperforms pure SFT algorithms.
Active Learning. The core concept behind active learning in machine learning theory is the ability of a model to interactively query a human or an oracle to obtain labels for new data points. Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative instances to query for labels. We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm. The literature has concentrated on label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy. For example, Cohn et al. (1994) demonstrates that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue. In the presence of noise, other papers focus on the classification problems (Hanneke 2007, Balcan et al. 2006, Hanneke 2011) and regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al. 2009), etc. Though our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
Data-driven Decision Making and Human in the Loop. A large body of recent literature on operations research or operations management studied how to effectively utilize data to improve their decisions, including pricing, assortment, recommendation, etc. Motivated by the observation that human may sometimes have “private” information which an algorithm does not have access to that can improve performance, Balakrishnan et al. (2022) examined new theory that describes algorithm overriding behavior. They designed and experimentally tested feature transparency as an implementable approach that can help people better identify and use their private information. Ibrahim et al. (2021) studied how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input. Dietvorst et al. (2018, 2015) study algorithm aversion phenomenon where people often fail to use algorithms after learning that they are imperfect, though evidence-based algorithms consistently outperform human forecasters. Other papers study Human-AI collaboration on a large variety of topics (Bastani et al. 2021, Wu et al. 2022, Deng et al. 2020), including ride-hailing (Benjaafar et al. 2022) and healthcare (Reverberi et al. 2022), etc.
Probabilistic Bisection. The other related literature, which is independent of the above human-AI ML topics, is probabilistic bisection algorithm (Horstein 1963). Algorithms utilizing the PBA framework have been employed in several contexts, notably including the task of target localization (Tsiligkaridis et al. 2014), scanning electron microscopy (Sznitman et al. 2013), topology estimation (Castro and Nowak 2008), edge detection (Golubev and Levit 2003), and value function approximation for optimal stopping problems (Rodriguez and Ludkovski 2015). From the computational perspective, Frazier et al. (2019) has shown that probabilistic bisection converges almost as quickly as stochastic approximation. In this work, we design the model alignment algorithm by utilizing the probabilistic bisection approach, where the randomness is involved due to the inherent variability in the human judgement process.
1.3 Organization of the Paper
@@
The remainder of this paper is organized as follows. Section 2 introduces the model setup. In Section 3, we introduce the model of human comparisons and characterize the sample complexity in a one-dimensional space. In Section 4, we present a two-stage framework which organically integrate SFT and human comparison. We specifically discuss the sample complexity when applying lasso in the first stage. Section 6 shows numerical experiments and managerial insights. Section 7 concludes the paper. Additional proofs and parameter calibrations are provided in the appendices.

### v2

@@
A Probabilistic Approach for Model Alignment with Human Comparisons
@@
A Probabilistic Approach for Model Alignment with Human Comparisons
@@
A growing trend involves integrating human knowledge into learning frameworks, leveraging subtle human feedback to refine AI models. While these approaches have shown promising results in practice, the theoretical understanding of when and why such approaches are effective remains limited. This work takes steps toward developing a theoretical framework for analyzing the conditions under which human comparisons can enhance the traditional supervised learning process. Specifically, this paper studies the effective use of noisy-labeled data and human comparison data to address challenges arising from noisy environment and high-dimensional models. We propose a two-stage “Supervised Learning+Learning from Human Feedback” (SL+LHF) framework that connects machine learning with human feedback through a probabilistic bisection approach. The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SL procedure and then uses human comparisons to improve the model alignment. To examine the efficacy of the alignment phase, we introduce a concept, termed the “label-noise-to-comparison-accuracy” (LNCA) ratio. This paper identifies from a theoretical perspective the conditions under which the “SL+LHF” framework outperforms the pure SL approach; we then leverage this LNCA ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity. We validate that the LNCA ratio meets the proposed conditions for its use through a case study conducted via Amazon Mechanical Turk (MTurk).
@@
model alignment, supervised learning, probabilistic bisection, human-AI interaction.
@@ 1 Introduction
In light of the increasing availability of data, the demand for data-driven algorithmic solutions has surged. Machine learning has emerged as the standard technique for numerous tasks, including natural language processing and computer vision, among others. Various streams of operations research also demonstrate its importance.
To model increasingly complex environments, machine learning models have grown in complexity, often leading to over-parameterization. While this increased complexity allows models to capture intricate patterns, it has revealed fundamental limitations in the standard training and testing pipeline. Specifically, the traditional supervised learning (SL) approach of training on noisy-labeled data faces significant challenges due to model misspecification (D’Amour et al. 2022). This issue becomes particularly acute in high-complexity scenarios, such as when the feature dimension is comparable to or exceeds the number of training data points. In such high-dimensional settings, fundamentally different models may achieve similar performance metrics during testing, despite having different underlying structures and behaviors. For instance, as demonstrated by Overman et al. (2024), the trained model that minimizes the empirical risk may violate fundamental principles, such as the natural expectation of negative demand elasticity—indicating that the model is not aligned with its expected behavior. Consequently, decision makers must consider strategic approaches to achieve better model alignment across different environments.
@@
One promising approach has been the incorporation of human feedback into the learning framework (Li et al. 2016, Christiano et al. 2017, Alur et al. 2024). These studies demonstrate that using human feedback for alignment—refining model parameters that are based on human-labeled data—can effectively address the limitations of traditional training approaches. For instance, in the design of chatbots, performance can be adaptively improved through the chatbots’ interactions with humans (Ziegler et al. 2019, Ouyang et al. 2022). Ouyang et al. (2022) exemplify this approach in their work on fine-tuning language models, where InstructGPT is fine-tuned with human feedback. The effectiveness of human feedback stems from evaluators’ domain expertise that extends beyond what can be encoded into a dataset or estimation method. Humans’ deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al. 2020).
Among various forms of human feedback, comparisons between alternatives have emerged as a particularly effective approach due to their cognitive simplicity. As Carl Jung observes, “Thinking is difficult, therefore let the herd pronounce judgment”—similarly, humans find it easier to select between alternatives than to generate absolute evaluations. Recent work has demonstrated the practical value of such comparative feedback: Zhou and Xu (2020) use pair-wise comparisons for story generation, while Xu et al. (2023) leverage human comparisons between large language models (LLMs) for fine-tuning. Although these applications show promising results, we lack a theoretical framework to understand when and why human comparisons enhance traditional supervised learning. This gap motivates our paper’s central question:
How can we effectively combine supervised learning and human feedback to achieve model alignment?
@@
To make progress towards answering this question, we consider a two-stage probabilistic model that helps us analyze and understand the interplay between supervised learning and human feedback. In the first stage, noisy labeled data are fed into the SL oracle to obtain low-dimensional representations (or embeddings). For example, selecting a small set of features through Lasso is a way to represent high-dimensional feature spaces; similarly, low-rank matrix estimation is an example of representation of high-dimensional matrices. The second stage uses human comparisons to improve model alignment within this low-dimensional space. Our framework uses the random utility model to describe humans’ comparison behavior. Practically, humans may provide incorrect responses, so we develop a probabilistic bisection approach for the alignment phase. We call this two-stage process the Supervised Learning+Learning from Human Feedback (SL+LHF) model.
The rationale behind this model is rooted in compelling evidence that representation learning techniques, such as deep learning architectures (Zhou and Paffenroth 2017, Akrami et al. 2019), can effectively learn low-dimensional features even from highly noisy data (Li et al. 2021, Taghanaki et al. 2021, Mousavi-Hosseini et al. 2024). For instance, Tsou et al. (2023) recently showed that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, sometimes even outperforming full-model fine-tuning. However, while supervised learning excels at learning these representations, the challenge lies in effectively fine-tuning the final layer—the stage where the model’s predictions are aligned with desired behaviors. For this alignment phase, we leverage human expertise through comparisons rather than direct labels, building on the cognitive advantage of comparative judgments discussed above. Our approach uses a sequential design of comparisons and active data acquisition to maximize information gain and minimize sample complexity, particularly effective when label noise is high.
@@
To describe the effectiveness of the refinement phase involving human comparison, we introduce a new concept called the “label-noise-to-comparison-accuracy” (LNCA) ratio. We offer a theoretical characterization of the condition in which the SL+LHF framework performs better than the pure SL algorithm based on this LNCA ratio. In a large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
@@ 1.1 Our Contributions
This paper makes contributions in four areas: theoretical modeling, characterization of sample complexity, practical implementation, and empirical validation.
Modeling.
We develop a theoretical framework that explains when and why human comparisons can substantially enhance performance of SL in high-noise environments. Our model captures the interaction between SL and human feedback through a utility-based comparative judgment approach.
Theoretical contributions.
Our main theoretical results consist of two components. First, we develop a probabilistic bisection method—an efficient sequential decision-making process that alternates between “vertical moves” (multiple evaluations of the same comparison) and “horizontal moves” (selection of new comparison pairs) to acquire human comparisons. Second, focusing on settings where the SL stage utilizes Lasso for learning low-dimensional representations, we introduce the label noise-to-comparison-accuracy ratio (LNCA) that quantifies the trade-off between label noise and human comparison accuracy. Our goal is to refine the model within an $\varepsilon$-distance with confidence at least $1-\delta$, which $\varepsilon$ represents the learning precision and $1-\delta$ denotes the confidence level. When human selection accuracy uniformly exceeds random guessing, we prove the refinement stage has sample complexity $O(s\log(1/(\delta\varepsilon)))$, where $s$ is the number of non-zero coefficients in Lasso. We characterize conditions under which our SL+LHF framework outperforms pure SL.
Towards practical implementations.
While our theoretical analysis examines human comparisons of model coefficients for analytical tractability, direct coefficient comparison is impractical in real applications. To bridge this gap, we develop an Active Sample Selection (ASS) framework that translates our theoretical insights into implementable procedures by enabling humans to compare meaningful data points rather than abstract model parameters. This framework provides a systematic approach for selecting the most informative samples for human comparison, applicable across a broad set of data distributions.
Empirical analysis.
Through Amazon Mechanical Turk (MTurk) experiments to acquire real human feedback on a stylized task, we validate our theoretical findings about the LNCA ratio and demonstrate the superior performance of SL+LHF over pure SL under conditions of higher observational noise, higher human comparison accuracy, and lower-dimensional feature spaces.
@@
The remainder of this paper is organized as follows: Section 2 reviews related literature. Section 3 presents our model setup, followed by Section 4, which introduces human comparison modeling and characterizes sample complexity in one dimension. Section 5 develops our two-stage framework integrating SL and human comparisons, with particular focus on sample complexity when using Lasso in the first stage. Section 6 details practical implementations for active sample selection, while Section 7 presents numerical experiments and managerial insights. Section 8 concludes. All proofs and parameter calibrations appear in the appendices.
2 Literature Review
Alignment and Learning from Human Feedback. The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms toward the objectives and interests of humans. Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts; this phenomenon has been identified as underspecification (D’Amour et al. 2022). Predictors that perform similarly in the training dataset can exhibit significant variation when they are deployed in other domains. The presence of ambiguity in a model might result in instability and suboptimal performance in real-world scenarios. Recent papers have focused on aligning models with human intentions, including training simple robots in simulated environments and Atari games (Christiano et al. 2017, Ibarz et al. 2018) and fine-tuning language models to summarize text (Stiennon et al. 2020) and to optimize dialogue (Jaques et al. 2019). Studies have demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al. 2022, Glaese et al. 2022, Stiennon et al. 2020, Dwaracherla et al. 2024). Ouyang et al. (2022) use RLHF (Stiennon et al. 2020) to fine-tune GPT-3 to follow a broad class of written instructions. In particular, Zhou and Xu (2020) use pair-wise comparisons for story generation. They find that, when comparing two natural language generation (NLG) models, asking a human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al. 2017), is more complicated. The much easier approach is for human annotators to directly compare one sample generated by the first model against another sample from the second model in a pair-wise fashion and then to compute the win/loss rate. Zhu et al. (2023) provide a reinforcement learning framework that includes human feedback – specifically, for function approximation in $K$-wise comparisons, with policy learning as the target. Xu et al. (2023) propose a fine-tuning algorithm by inducing a complementary distribution over the two sampled LLM responses to model human comparison processes, in accordance with the Bradley-Terry model for pairwise comparisons. In contrast to all these works, we provide a theoretical, two-stage framework for integrating supervised learning with actively acquired human comparison labels, and we characterize the conditions where the two-stage framework outperforms pure SL algorithms.
Transfer Learning and Task Alignment. Model alignment generally refers to the process of ensuring that machine learning models are tailored to meet specific tasks and objectives. A related paradigm is transfer learning, where a model trained on one task is adapted for use in another task through additional fine-tuning (Pan and Yang 2009, Xu et al. 2021, Bastani et al. 2022, Du et al. 2024). While our work shares some conceptual similarities with transfer learning, we focus specifically on the impact of noise in both supervised learning and human feedback phases. Our framework’s applicability to transfer learning scenarios depends crucially on the relationship between the initial and target tasks’ representations. When the target task shares the same low-dimensional representation as the initial task, requiring only updates to the final layer or coefficients, our theoretical results apply directly—the human comparison phase can effectively refine these parameters regardless of whether the comparison objective exactly matches the initial task. However, when the target task requires different representations, transfer learning typically demands an additional supervised learning phase with new labeled data to first refine these representations before any human-guided fine-tuning can begin. In such cases, our framework’s human comparison stage remains relevant for the final alignment when label noise is high, but it must be preceded by this representation learning phase.
Active Learning. The core concept behind active learning in machine learning theory is the ability of a model to interactively query a human or an oracle to obtain labels for new data points. Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative learning instances to query for labels. We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm literature. The literature has concentrated on studying label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy. For example, Cohn et al. (1994) demonstrate that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue. In the presence of noise, other papers focus on the classification problems (Hanneke 2007, Balcan et al. 2006, Hanneke 2011) and the regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al. 2009). Although our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
Data-Driven Decision Making and Humans in the Loop. A large body of recent literature on operations research or operations management has studied how to effectively use data to improve decisions (Mišić and Perakis 2020), including pricing, assortment, and recommendation. Motivated by the observation that humans may sometimes have “private” information to which an algorithm does not have access but that can improve performance, Balakrishnan et al. (2022) examine new theory that describes algorithm overriding behavior. They designed an experiment to test feature transparency as an implementable approach that can help people better identify and use their private information. Ibrahim et al. (2021) study how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input. Dietvorst et al. (2018, 2015) study the algorithm aversion phenomenon, in which people fail to use algorithms after learning that they are imperfect, even though evidence-based algorithms consistently outperform human forecasters. Other papers study human-artificial intelligence (AI) collaboration on a large variety of topics (Bastani et al. 2021, Wu et al. 2022, Deng et al. 2020, Chen et al. 2022, Caro and de Tejada Cuenca 2023, Ye et al. 2024), including ride-hailing (Benjaafar et al. 2022) and health care (Reverberi et al. 2022).
Probabilistic Bisection. Another stream of literature related to our work, which is independent of the previously identified human-AI ML topics, is on probabilistic bisection algorithms (PBAs) (Horstein 1963). Algorithms using the PBA framework have been used in several contexts – notably, including the tasks of target localization (Tsiligkaridis et al. 2014), scanning electron microscopy (Sznitman et al. 2013), topology estimation (Castro and Nowak 2008), edge detection (Golubev and Levit 2003), and value function approximation for optimal stopping problems (Rodriguez and Ludkovski 2015). From a computational perspective, Frazier et al. (2019) have shown that probabilistic bisection converges almost as quickly as stochastic approximation. In this work, we design the model alignment algorithm by using the PBA approach, where randomness is involved because of the inherent variability in the human judgment process.